# Universality of Efimov states in highly mass-imbalanced cold-atom mixtures with van der Waals and dipole interactions

### v2

@@ Acknowledgements.
We thank Yoshiro Takahashi, Yousuke Takasu, Lucas Happ, and Emiko Hiyama for fruitful discussions. This work was supported by JSPS KAKENHI Grant Numbers JP22K03492, JP23H01174, and JP23K03292. KO acknowledges support from Graduate Program on Physics for the Universe (GP-PU) of Tohoku University. SE acknowledges support from Institute for Advanced Science, University of Electro-Communications.
@@
where $\nu_{0}=s_{\ell}/2$. In the absence of the induced attraction via the light particle, Eq. (2) is the van der Waals potential problem; the attraction modifies the last term in the left-hand side from the quarter integer $\nu_{0}=(2\ell+1)/4$ into $\nu_{0}=s_{\ell}/2$. When the Efimov effect is present (absent) as in the $\ell=0$ (large $\ell$) channel, $\nu_{0}$ is pure imaginary (real).
In a similar manner as Ref. [51], the two linearly independent solutions of Eq. (2) are repretented as
@@
$b_{n}$ and $\nu$ are determined from the same trandental equations in Ref. [51], with the modification $\nu_{0}=(2\ell+1)/4\rightarrow\nu_{0}=s_{\ell}/2$. In order to have favorable short-range asymptotc features, it is conveniet to consider $f_{\ell}^{c}$ and $g_{\ell}^{c}$ which are defined as linear transforms of $\overline{f_{\ell}}$ and $\overline{g_{\ell}}$ as
@@
$\begin{split}&f^{c}_{\ell}(r)\rightarrow\frac{1}{\sqrt{2\pi}}\frac{r^{3/2}}{r_{\rm vdw}}\cos\left[x-\frac{\pi}{4}\right]\\ &g^{c}_{\ell}(r)\rightarrow-\frac{1}{\sqrt{2\pi}}\frac{r^{3/2}}{r_{\rm vdw}}\sin\left[x-\frac{\pi}{4}\right].\end{split}$ (25)
Notably, they are independent of angular momentum nor energy, which are essential features for the angular-insensitive QDT [52].
@@ Appendix B Low-energy QDT expansion
$\nu=\nu_{0}+\frac{3}{2\nu_{0}}\frac{\Delta^{2}}{(\nu_{0}^{2}-1)(4\nu_{0}^{2}-1)}+O(\Delta^{4}),$ (30)
@@
Notably, all the above equations remain unchanged from Refs. [53, 54] even when $\nu$ is real non-quater integer value ($s_{\ell}^{2}>0$), or a pure imaginary value ($s_{\ell}^{2}<0$).
@@
On the other hand, in the presence of the Efimov effect ($\nu_{0}=s_{\ell}/2$ with $s_{\ell}^{2}<0$), $\nu$ takes a pure imaginary value. Using
@@
As the angular momentum is the good quantum number when $\hat{V}_{dd}=0$, we label the states with $L$ and $I$ as |$\ket{\psi_{IL}}\simeq\ket{\psi^{(0)}_{IL}}$ : $L$ is the angular momentum of the unperturbed state and $I$ is the label of the state. The second order perturbation theory reads
$E_{IL}$ $=E^{(0)}_{IL}+E^{(1)}_{IL}+E^{(2)}_{IL}$ (42)
@@
$E^{(1)}_{IL}$ $=\bra{\psi^{(0)}_{IL}}\hat{V}_{dd}\ket{\psi^{(0)}_{IL}}$ (44)
$E^{(2)}_{IL}$ $=-\sum_{i\ell\neq IL}\frac{\left|\bra{\psi^{(0)}_{IL}}\hat{V}_{dd}\ket{\psi^{(0)}_{i\ell}}\right|^{2}}{E^{(0)}_{i\ell}-E^{(0)}_{IL}}$ (45)
$a^{(1)}_{{IL},i\ell}$ $=-\frac{\bra{\psi^{(0)}_{i\ell}}\hat{V}_{dd}\ket{\psi^{(0)}_{IL}}}{E^{(0)}_{i\ell}-E^{(0)}_{IL}}$ (46)
@@
where the bracket in the last line is the Wigner’s $3j$ symbol. We note that the azimuthal number $m$ around the $z$ axis is conserved even in the presence of the dipole interaction and thus unchanged. As the $3j$ symbol only takes non-zero values for $\ell=L,L\pm 2$, the perturbation only couples $+2,0,-2$ angular momentum states. Introducing the matrix element of the radial part
@@
$\begin{split}E_{I,L=0}&=E^{(0)}_{I,L=0}-\frac{4}{5}\left(\frac{3a_{dd}}{M}\right)^{2}\sum_{i}\frac{\left|\langle i,\ell=2|v_{r}|I,L=0\rangle\right|^{2}}{E_{i,\ell=2}^{(0)}-E_{I,L=0}^{(0)}}\end{split}$ (51)
@@
$E_{I,L=1}=$ $E^{(0)}_{I,L=1}-\frac{4}{5}\frac{3a_{dd}}{M}\langle I,L=1|v_{r}|I,L=1\rangle$ (53)
@@
Figure 11: Binding energy of the bosonic dimers of two heavy atoms (green triangles) and trimers (blue circles). The physical parameter in each row is the same as Fig. 5. The left panels are plotted against $a_{\mathrm{no-dd}}$ while the right panels are against $a_{\mathrm{with-dd}}^{(\mathrm{HH})}$, both in their positive scattering length side. The others are the same as Fig. 5.
@@ Appendix D Dimer energy spectra
For a moderate dipole interaction strength $a_{dd}\sim r_{\mathrm{vdw}}$. the dimer energies for the bosonic system are shown as green points in Fig. 11. While the dimer energies plotted against $a_{\mathrm{no-dd}}^{(\mathrm{HH})}$ seem to differ significantly from those in the pure van der Waals system (green dashed curve in Fig. 2(b), they can be converted in a more universal form by introducing the $s$-wave scattering length incorporating the channel couplings $a_{\mathrm{with-dd}}^{(\mathrm{HH})}$, in the same manner as Sec. III.3. This universal behavior of the two-body dipolar scattering is consistent with Refs. [59, 60]: the energy of the dimer of $s$-wave dominant character is universally described by the $s$-wave scattering length $E\propto a^{-2}$. There also appear dimers of higher angular momentum, adjacent to the $s$-wave one; they behave universally with respect to different values of $R_{\mathrm{min}}$, but not as much as compared to the $s$-wave one. We have found that the spreading of the data points decrease as $R_{\mathrm{min}}$ is decreased, suggesting that this higher-angular-momentum dimer gradually converges to a universal behavior, with a slowness of convergence similar to that of the trimers in the avoided crossing region. We also show the trimer energies (the same data points in Fig. 5) for comparison. As in Fig. 2(b), some of the trimers’ energies lie above those of the dimers, so that the tirmers are embedded in the dimer+atom continuum. We note that the trimers appear as bound states in our Born-Oppenheimer calculation because the continuum channels are neglected. There are two notable dimer curves; one of them is the $s$-wave dominant dimer, which behaves in the same manner as that in Fig. 11 when plotted via $a_{\mathrm{with-dd}}^{(\mathrm{HH})}$. The other one is of $d$-wave dominant nature, which appears at around $a_{\mathrm{no-dd}}^{(\mathrm{HH})}/r_{\mathrm{vdw}}\simeq 1$, and seems to lie more or less besides the trimers in the avoided crossing region if plotted via $a_{\mathrm{with-dd}}^{(\mathrm{HH})}$.
@@
In Fig. 12. dimer energies for the fermionic system are shown, for the parameters corresponding to Fig. 8. As it is impossible to introduce the $p$-wave scattering volume, it is not easy to present it in an analogous way as Fig. 2(d). Still, most of the dimer energies show universal behaviors with respect to $K_{c,F}^{(3D)}$, together with some other stray points which we suspect to be of high-angular-momentum character $\ell\geq 3$. Similarly to Fig. 2(d), some of the trimers lie above the dimer energy, and are thus embedded in the dimer-atom continuum.

### v3

@@ Acknowledgements.
We thank Yoshiro Takahashi, Yousuke Takasu, Lucas Happ, and Emiko Hiyama for fruitful discussions. This work was supported by JSPS KAKENHI Grants No. JP22K03492, JP23H01174, and JP23K03292. K.O. acknowledges support from Graduate Program on Physics for the Universe (GP-PU) of Tohoku University. S.E. acknowledges support from the Institute for Advanced Science, University of Electro-Communications.
@@
where $\nu_{0}=s_{\ell}/2$. In the absence of the induced attraction via the light particle, Eq. (2) is the van der Waals potential problem; the attraction modifies the last term in the left-hand side from the quarter integer $\nu_{0}=(2\ell+1)/4$ into $\nu_{0}=s_{\ell}/2$. When the Efimov effect is present (absent) as in the $\ell=0$ (large $\ell$) channel, $\nu_{0}$ is purely imaginary (real).
In a similar manner as Ref. [51], the two linearly independent solutions of Eq. (2) are represented as
@@
$b_{n}$ and $\nu$ are determined from the same transcendental equations in Ref. [51], with the modification $\nu_{0}=(2\ell+1)/4\rightarrow\nu_{0}=s_{\ell}/2$. In order to have favorable short-range asymptotic features, it is convenient to consider $f_{\ell}^{c}$ and $g_{\ell}^{c}$, which are defined as linear transforms of $\overline{f_{\ell}}$ and $\overline{g_{\ell}}$ as
@@
$\begin{split}&f^{c}_{\ell}(r)\rightarrow\frac{1}{\sqrt{2\pi}}\frac{r^{3/2}}{r_{\rm vdw}}\cos\left[x-\frac{\pi}{4}\right],\\ &g^{c}_{\ell}(r)\rightarrow-\frac{1}{\sqrt{2\pi}}\frac{r^{3/2}}{r_{\rm vdw}}\sin\left[x-\frac{\pi}{4}\right].\end{split}$ (25)
Notably, they are independent of angular momentum nor energy, which are essential features for the angular-momentum-insensitive QDT [52].
@@ Appendix B Low-energy QDT expansion
$\nu=\nu_{0}-\frac{3}{2\nu_{0}}\frac{\Delta^{2}}{(\nu_{0}^{2}-1)(4\nu_{0}^{2}-1)}+O(\Delta^{4}),$ (30)
@@
Notably, all the above equations remain unchanged from Refs. [53, 54] even when $\nu$ is a real non-quater integer value ($s_{\ell}^{2}>0$), or a purely imaginary value ($s_{\ell}^{2}<0$).
@@
On the other hand, in the presence of the Efimov effect ($\nu_{0}=s_{\ell}/2$ with $s_{\ell}^{2}<0$), $\nu$ takes a purely imaginary value. Using
@@
As the angular momentum is the good quantum number when $\hat{V}_{dd}=0$, we label the states with $L$ and $I$ as |$\ket{\psi_{IL}}\simeq\ket{\psi^{(0)}_{IL}}$ : $L$ is the angular momentum of the unperturbed state and $I$ is the label of the state. The second-order perturbation theory reads
$E_{IL}$ $=E^{(0)}_{IL}+E^{(1)}_{IL}+E^{(2)}_{IL},$ (42)
@@
$E^{(1)}_{IL}$ $=\bra{\psi^{(0)}_{IL}}\hat{V}_{dd}\ket{\psi^{(0)}_{IL}},$ (44)
$E^{(2)}_{IL}$ $=-\sum_{i\ell\neq IL}\frac{\left|\bra{\psi^{(0)}_{IL}}\hat{V}_{dd}\ket{\psi^{(0)}_{i\ell}}\right|^{2}}{E^{(0)}_{i\ell}-E^{(0)}_{IL}},$ (45)
$a^{(1)}_{{IL},i\ell}$ $=-\frac{\bra{\psi^{(0)}_{i\ell}}\hat{V}_{dd}\ket{\psi^{(0)}_{IL}}}{E^{(0)}_{i\ell}-E^{(0)}_{IL}},$ (46)
@@
where the bracket in the last line is the Wigner’s $3j$ symbol. We note that the azimuthal number $m$ around the $z$ axis is conserved even in the presence of the dipole interaction and thus unchanged. As the $3j$ symbol only takes non-zero values for $\ell=L,L\pm 2$, the perturbation couples only $+2,0,-2$ angular momentum states. Introducing the matrix element of the radial part
@@
$\begin{split}E_{I,L=0}&=E^{(0)}_{I,L=0}-\frac{4}{5}\left(\frac{3a_{dd}}{M}\right)^{2}\sum_{i}\frac{\left|\langle i,\ell=2|v_{r}|I,L=0\rangle\right|^{2}}{E_{i,\ell=2}^{(0)}-E_{I,L=0}^{(0)}},\end{split}$ (51)
@@
$E_{I,L=1}=$ $E^{(0)}_{I,L=1}-\frac{4}{5}\frac{3a_{dd}}{M}\langle I,L=1|v_{r}|I,L=1\rangle,$ (53)
@@
Figure 11: Binding energy of the bosonic dimers of two heavy atoms (green triangles) and trimers (blue circles). The physical parameter in each row is the same as Fig. 5. The left panels are plotted against $a_{\mathrm{no-dd}}$, and the right panels are against $a_{\mathrm{with-dd}}^{(\mathrm{HH})}$, both in their positive scattering length side. The others are the same as Fig. 5.
@@ Appendix D Dimer energy spectra
For a moderate dipole interaction strength $a_{dd}\sim r_{\mathrm{vdw}}$. the dimer energies for the bosonic system are shown as green points in Fig. 11. While the dimer energies plotted against $a_{\mathrm{no-dd}}^{(\mathrm{HH})}$ seem to differ significantly from those in the pure van der Waals system (green dashed curve in Fig. 2(b), they can be converted into a more universal form by introducing the $s$-wave scattering length incorporating the channel couplings $a_{\mathrm{with-dd}}^{(\mathrm{HH})}$, in the same manner as Sec. III.3. This universal behavior of the two-body dipolar scattering is consistent with Refs. [59, 60]: the energy of the dimer of $s$-wave dominant character is universally described by the $s$-wave scattering length $E\propto a^{-2}$. There also appear dimers of higher angular momentum, adjacent to the $s$-wave one; they behave universally with respect to different values of $R_{\mathrm{min}}$, but not as much as compared to the $s$-wave one. We have found that the spreading of the data points decrease as $R_{\mathrm{min}}$ is decreased, suggesting that this higher-angular-momentum dimer gradually converges to a universal behavior, with a slowness of convergence similar to that of the trimers in the avoided crossing region. We also show the trimer energies (the same data points in Fig. 5) for comparison. As in Fig. 2(b), some of the trimers’ energies lie above those of the dimers, so that the trimers are embedded in the dimer+atom continuum. We note that the trimers appear as bound states in our Born-Oppenheimer calculation because the continuum channels are neglected. There are two notable dimer curves; one of them is the $s$-wave dominant dimer, which behaves in the same manner as that in Fig. 11 when plotted via $a_{\mathrm{with-dd}}^{(\mathrm{HH})}$. The other one is of $d$-wave dominant nature, which appears at around $a_{\mathrm{no-dd}}^{(\mathrm{HH})}/r_{\mathrm{vdw}}\simeq 1$, and seems to lie more or less besides the trimers in the avoided crossing region if plotted via $a_{\mathrm{with-dd}}^{(\mathrm{HH})}$.
@@
In Fig. 12. dimer energies for the fermionic system are shown, for the parameters corresponding to Fig. 8. As it is impossible to introduce the $p$-wave scattering volume, it is not easy to present it in an analogous way to Fig. 2(d). Still, most of the dimer energies show universal behaviors with respect to $K_{c,F}^{(3D)}$, together with some other stray points which we suspect to be of high-angular-momentum character $\ell\geq 3$. Similarly to Fig. 2(d), some of the trimers lie above the dimer energy, and are thus embedded in the dimer-atom continuum.